If you have the digits 0-6 and there is no restrictions on repetitions and/or zeros, calculate how many possible 4 digit arrangements you can have that contain at least one even digit. Show work
Answer #1

Distinct Numbers with Repetition, All Numbers are used: If an n distinct number string has to be arranged, with repetition using all the numbers say x, then the total number of ways of arranging the same is nx.

The digits are 0, 1, 2 , 3, 4, 5 and 6...7 digits of which 2, 4 and 6 (3 digits) are even and the rest 0, 1, 3 and 5 (4 digits) are not even.

The total number of arrangements = 7 * 7 * 7 * 7 = 2401

Number of arrangements where there is not a single even number = 4 * 4 * 4 * 4 = 256

Therefore number of arrangements which have at least 1 even number = 2401 - 256 = 2145